### v2

# Skip-Vision: Efficient and Scalable Acceleration of Vision-Language Models via Adaptive Token Skipping

Abstract
Transformer-based models have driven significant advancements in Multimodal Large Language Models (MLLMs), yet their computational costs surge drastically when scaling resolution, training data, and model parameters. A key bottleneck stems from the proliferation of visual tokens required for fine-grained image understanding. We propose Skip-Vision, a unified framework addressing both training and inference inefficiencies in vision-language models. On top of conventional token compression approaches, our method introduces two complementary acceleration strategies. For training acceleration, we observe that Feed-Forward Network (FFN) computations on visual tokens induce marginal feature updates. This motivates our Skip-FFN strategy, which bypasses FFN layers for redundant visual tokens. For inference acceleration, we design a selective KV-cache removal mechanism that prunes the skipped key-value pairs during decoding while preserving model performance. Experimental results demonstrate that Skip-Vision reduces training time by up to 35%, inference FLOPs by 75%, and latency by 45%, while achieving comparable or superior performance to existing methods. Our work provides a practical solution for scaling high-performance MLLMs with enhanced efficiency.
1 Introduction
Transformer-based large-scale models have significantly advanced progress in Artificial Intelligence (AI) [87, 10, 102]. Their remarkable capabilities have driven the emergence of modern Multimodal Large Language Models (MLLMs) [3, 69, 89, 2], which demonstrate vision-language competencies comparable to human performance. Empirical evidence suggests that scaling laws continue to be effective, primarily across three dimensions: visual scaling [67, 112, 40], data scaling [64, 63], and model scaling [105, 56]. However, these scaling methods present substantial efficiency issues, significantly increasing the training and inference burden. In a typical LLaVA-1.5 framework [69], where the model receives 576 visual tokens and 40 text tokens, inference using Vicuna-13B [24] LLM backbone demands 200 trillion FLOPS [42]. These motivate our pursuit of an MLLM architecture that optimally balances efficiency and performance.
Figure 1: Performance-efficiency trade-off curve. Each circle denotes a model configuration, where our models utilize the Skip-Vision framework, with CoS [40], LLava-HD [67] and LLaVA [69] serving as baselines. Circle sizes reflect the inference FLOPs ratio. Skip-Vision demonstrate superior performance, scaling effectively with increased FLOPs and data and achieving higher inference efficiency when compared to baselines and other effcient MLLM methods. All methods utilize LLaMA3 8B as the foundational large language model.
Recent advances in Multimodal Large Language Models (MLLMs), such as Chain-of-Sight (CoS)[40] and LLaVA-Next[68], have improved performance by increasing visual tokens. However, during the Supervised Fine-Tuning (SFT) stage, the computational cost rises dramatically, leading to diminishing returns and making the scaling unsustainable. As shown in Figure 1, CoS requires a $165\%$ increase in training resources and a $400\%$ increase in inference computation for a mere $1.7\%$ performance improvement, while LLaVA-Next faces similar inefficiencies. This challenge emphasizes the need to reduce visual token computation. Approaches like token compression or pruning [60, 116, 21, 39, 119, 109], while efficient, risk losing critical visual information, revealing a trade-off between efficiency and performance. This work aims to optimize the model’s computational efficiency from an architectural perspective, proposing a scalable solution that balances efficiency with visual fidelity.
Llama layer Complexity Parameters
FFN $10.5LNC^{2}$ $5.64B(70\%)$
Attention $2LN^{2}C+1.5NC^{2}$ $1.34B(17\%)$
Embedding - $0.5B(6\%)$
Table 1: Computational complexity and parameter statistics for Llama3 8B. L denotes the number of blocks, N represents the number of input tokens, and C is the feature dimensionality
We systematically examine the parameter count and FLOPs distribution across different components of the LLM. As shown in Table 1, using LLaMA-3 [28] as an example, the parameter and FLOPS requirements for the FFN component are significantly higher than those for the attention module. Furthermore, as illustrated in Figure 2, when the number of visual tokens is large, the majority of computations are concentrated on FFN of these tokens. In particular, we computed the difference in the magnitudes of the features before and after the FFN transformation for different tokens. As illustrated in the figure 3, we observe that the updates introduced by the FFN to the visual tokens are significantly smaller than those that affect the text tokens. This observation suggests a potential approach: could we bypass FFN computations for a substantial portion of visual tokens? Furthermore, in the autoregressive training framework of MLLMs, only the final visual token utilizes its last hidden state for next-token prediction, while the other visual tokens primarily serve to convey causal information. During model inference, inspired by [7], we observe that most visual information is progressively integrated into the final token through the early causal layers. This raises the possibility of reducing reliance on the full visual scaling KV cache throughout inference, potentially accelerating the inference process.
Figure 2: Flops ratio of visual tokens in FFN. When visual tokens greatly outnumber text tokens, computation is predominantly consumed by the FFN of visual tokens. Reducing their passage through the FFN can thus markedly decrease computational costs.
Figure 3: FFN imapct for different MLLMs. We evaluate FFN impact by computing the feature modulus ratio before ($\|h_{\text{attn}}\|_{2}$) and after FFN ($\|\text{FFN}(h_{\text{attn}})\|_{2}$). The FFN introduces significantly smaller updates to visual tokens compared to text tokens.
Building on these insights, we propose Skip-Vision, a novel and flexible architecture designed for more efficient scaling (see Figure 1 for a comparison). Visual tokens derived from smaller window/scale sizes constitute approximately $80\%$ of the total visual tokens and account for most of the computational overhead. To mitigate this, as shown in Figure 4, we allow these tokens to skip the Feed-Forward Network (FFN) computation, significantly reducing overall computational costs. To address redundancy within skipped vision tokens, we can apply a token merging strategy for further simplification. During inference, skipped tokens are removed from the KV cache after the pre-fill phase, accelerating subsequent computation. Empirical results demonstrate that Skip-Vision reduces training time by up to $35\%$, FLOPs of inference by $75\%$ and inference latency by $45\%$, while maintaining or even improving performance. These efficiency gains provide a scalable solution for large-scale multimodal learning, enabling better data utilization and enhanced model scalability.
In summary, our contributions are as follows.
  • We introduce a novel and efficient Skip-Vision framework, using token merge and skip FFN strategy during training to reduce redundant computation for visual tokens.
  • In inference, our framework employs a skip KV-cache mechanism that removes skip-FFN visual tokens from the KV-cache, enhancing efficiency.
  • We present a theoretical analysis of the performance error in Skip-Vision and establish an error bound, which will be empirically validated through experiments.
  • Experiments show our model’s superior efficiency, effective data scaling, and performance on par with state-of-the-art models of similar scale.
2 Related work
2.1 Vison-Language models
Rapid progress in large language models (LLMs) [28, 24, 114] has laid a solid foundation for the emergence of large-scale vision-language models (VLMs) [3, 58, 6, 18]. Flamingo [3] was a pioneering effort, integrating pre-trained image encoders with LLMs to handle multimodal data, thereby allowing comprehensive understanding and reasoning across modalities. Following this, the major models developed by leading corporations, such as GPT-4v [2] and Gemini-1.5-Pro [89], have led the progress in this domain. Using proprietary data sets and undisclosed training methodologies, these models have elevated multimodal intelligence to unprecedented levels.
At the same time, the open-source community has worked diligently to stay abreast of these advancements. The LLava series models [69, 67, 112, 56] exemplify a key strategy in this domain by mapping visual features directly into the input embedding space of the language model, integrating them as input tokens. However, this approach generates a substantial volume of visual tokens, frequently exceeding the number of text tokens, creating efficiency challenges. InternVL-1.5 [23] adopts dynamic high-resolution techniques, segmenting large images into smaller fragments for processing and scaling them accordingly. MiniCPM-V [117] utilizes a perceiver resampler structure with a single layer of cross-attention to compress visual tokens. Models such as Vary [108], SPHINX [64], Cambrian-1 [101], and BRAVE [47] leverage multi-visual encoder architectures to strengthen their visual processing capabilities. Furthermore, a crucial factor contributing to the effectiveness of modern VLMs is instruction fine-tuning [27, 131, 67], enabling these models to function as conversational agents and interact with users through natural, human-like dialogue.
2.2 Efficient multimodal large language models
To enhance the efficiency of vision-language models, one of the primary and widely employed approaches in contemporary research involves reducing the size of the backbone models [26, 123, 130, 36, 133, 83]. Simultaneously, given that visual tokens contribute significantly to computational demands, research focused on compressing or refining these tokens has garnered increasing attention. Such methods are often coupled with the design of vision-language projectors. For instance, techniques like the Perceiver Resampler [41, 6, 3] and Q-Former [58] utilize transformer-based mechanisms to consolidate visual tokens into more compact query sets. In this way, the transformer outputs corresponding to the positions of the learnable latent queries serve as the aggregated representation of the visual features [42]. Honeybee [14] propose two visual projectors, namely C-Abstractor and D-Abstractor. LLaVA-PruMerge [92] and MADTP [12] introduce adaptive methods for reducing visual tokens. In addition, while the large language model (LLM) field has employed various techniques for token reduction to accelerate inference and compress KV cache [34, 128], research in this area remains limited within the MLLM domain.
3 Preliminaries
Figure 4: The framework of Skip-Vision. a) While visual scaling enriches visual information, it also increases computational overhead. Skip-Vision uses a skip-FFN strategy during training to reduce redundant computation for visual tokens. The numerous skipped tokens will be limited to the attention layer and bypass FFN. b) At the beginning of inference, Skip-Vision will remove skip-FFN visual tokens from the initial KV-cache, enhancing efficiency. c) During inference, skip attention leverages the skip KV-cache to accelerate generation.
3.1 Decoder-only LLM
Each Transformer layer in a decoder-only LLM primarily consists of a self-attention layer and a feed-forward neural network. To facilitate autoregressive inference, the use of a KV-cache has been developed to accelerate the decoding process.
Self-Attention Mechanism: The self-attention mechanism enables each token in a sequence to selectively attend to other tokens, capturing contextual dependencies. Each token generates a query $(Q)$, key $(K)$, and value $(V)$ vector. Attention scores are calculated by the scaled dot-product:
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
where $d_{k}$ is the dimension of $K$. This formulation allows each token to incorporate information from others based on relevance, making self-attention critical in capturing relationships across tokens.
Feed-Forward Neural Network: The feed-forward neural network enhances token representations following the self-attention layer. The operation can be expressed as:
$\operatorname{FFN}(x)=\operatorname{Activation}\left(xW_{1}+b_{1}\right)W_{2}+b_{2},$ (2)
where $x$ is the input, $W_{1}$ and $W_{2}$ are weight matrices, and $b_{1}$ and $b_{2}$ are biases. This architecture allows the model to capture complex patterns and improve the expressiveness of token representations.
KV-cache: The key-value cache (KV-cache) stores key $(K)$ and value $(V)$ representations from previous self-attention computations, enabling efficient reuse during inference. This approach avoids redundant calculations for earlier tokens, improving inference speed.
The attention score is computed using $\operatorname{Attention}\left(Q,K^{t}_{\text{cache }},V^{t}_{\text{cache }}\right)$, where $Q$ is the query for the current token. At each time step $t$, the cache is updated as:
$K_{\text{cache }}^{(t)}=\left[K_{\text{cache }}^{(t-1)};K_{t}\right],\quad V_{\text{cache }}^{(t)}=\left[V_{\text{cache }}^{(t-1)};V_{t}\right],$ (3)
where $K_{t}$ and $V_{t}$ are the current token’s key and value. This mechanism enhances computational efficiency and accelerates token generation.
4 Method
4.1 Skipped token selection and token merge
Before applying the skip FFN strategy, we first identify the skipped and retained visual tokens. The retained tokens should preserve most essential visual information, while the skipped tokens serve as complementary details, ensuring efficiency without significant loss of information.
Fortunately, in MLLM architectures, we can naturally distinguish retained and skipped tokens based on global and local context tokens. Global context tokens are retained, while local context tokens are skipped. In architectures like LLaVA-HD, global tokens originate from the resized full image, whereas local tokens come from smaller image patches. In CoS, global tokens correspond to larger window sizes, while local tokens are derived from smaller windows. Since LLaVA lacks a clear distinction between global and local tokens, inspired by [115], we select the top-$n$ tokens most similar to the CLS token as retained tokens, categorizing the rest as skipped tokens.
To further reduce the computational burden, it is necessary to refine the visual tokens to eliminate excessive redundancy. As shown in Figure 5, we conducted a similarity analysis of the visual tokens in CoS at different levels, calculating their similarity density. We found that the smaller the granularity of the visual tokens, the higher their average similarity, while their token count is significantly greater than that of other granularities. Consequently, inspired by ToMe [9], we employ a token merge method to reduce redundancy in skipped tokens, as detailed below:
  • Compute cosine similarities: calculate the cosine similarity between each pair of tokens in the set, obtaining an average similarity score for each token.
  • Select top $k$ distinctive Tokens: sort tokens by average similarity in ascending order and retain the top $k$ tokens with the lowest average similarity, representing the most distinctive information.
  • Merge remaining tokens: for the remaining $n-k$ tokens, merge each one with the most similar token among the top $k$, thus consolidating information efficiently.
4.2 Training speed-up: skip FFN
As depicted in Figure 4a, we propose the FFN skip strategy: The retained visual tokens, which are few in number, proceed conventionally through all decoder layers of the LLM, while the numerous skipped visual tokens will be limited to the self-attention layers of each transformer block. For a given layer $l$, the output of retained tokens $h_{\text{retained}}^{(l)}$ is the sum of the self-attention output $h_{\text{attn}}^{(l)}$ and the FFN-transformed features $\text{FFN}^{(l)}(h_{\text{attn}}^{(l)})$:
$h_{\text{retained}}^{(l)}=h_{\text{attn}}^{(l)}+\text{FFN}^{(l)}(h_{\text{attn}}^{(l)}).$ (4)
The output of the skipped tokens is:
$h_{\text{skipped}}^{(l)}=h_{\text{attn}}^{(l)}.$ (5)
However, bypassing the FFN layers presents a non-trivial challenge. In the autoregressive training process, the final token of the vision sequence is used to predict the initial token of the text sequence, necessitating a specialized design for this terminal vision token. To facilitate this transition effectively, it must pass through the FFN layers, which are enriched with stored linguistic information [84, 31].
To address this, as shown in Figure 4, we introduce adaptive summary tokens to efficiently consolidate visual information. This token is generated by the Adaptive Summary Layer before the token sequence is input into the LLM. The layer comprises a simple linear transformation, and its computation is defined as follows:
$x^{s}=(\operatorname{softmax}(X\cdot W^{T}))^{T}\cdot X,$ (6)
where $X$ denotes the retained or skipped visual tokens, $x^{s}\in\mathbb{R}^{1\times C}$ is the summary token, and $W\in\mathbb{R}^{1\times C}$ is the weight matrix. Leveraging this layer, we concatenate a summary token on both ends of the skipped visual tokens before their input to the LLM. The first summary token aggregates essential information from retained tokens, mitigating potential loss from extended token sequences. The second summary token integrates prior sequence information and plays a key role in predicting the next text token.
Figure 5: Similarity density statistics. We conducted a similarity density analysis in CoS, examining the visual tokens derived from various window scale. The results indicate that as the window scale decreases, the similarity between visual tokens becomes more pronounced, reflecting higher token redundancy.
4.3 Inference speed-up: skip KV-cache
After training, our Skip-Vision framework incorporates skip KV-cache method to further improve efficiency:
$K_{\text{skip-cache }}^{(0)}=K_{\text{cache }}^{(0)}\setminus K_{skip},\quad V_{\text{skip-cache }}^{(0)}=V_{\text{cache }}^{(0)}\setminus V_{skip},$ (7)
As depicted in Figures 4 b) and c), given the autoregressive nature of LLMs, earlier token information is progressively transferred to later tokens. Skip FFN further enhances this integration by suppressing the attention weights of skipped tokens, forcing the model to focus on key tokens and improving information aggregation efficiency. Empirical results show that, due to the inclusion of summary tokens, this omission does not compromise model performance, underscoring a distinctive advantage of our framework for more efficient inference.
5 Theoretical analysis of Skip-Vision
At the core of this analysis lies the layer error incurred when bypassing the FFN layer. For a given layer $l$, the original output $h_{\text{original}}^{(l)}$ is : $h_{\text{original}}^{(l)}=h_{\text{attn}}^{(l)}+\text{FFN}^{(l)}(h_{\text{attn}}^{(l)}).$ When skipping the FFN, the output becomes: $h_{\text{skip}}^{(l)}=h_{\text{attn}}^{(l)}.$ The per-layer skipping error is:
$\epsilon^{(l)}=\|h_{\text{original}}^{(l)}-h_{\text{skip}}^{(l)}\|_{2}=\|\text{FFN}^{(l)}(h_{\text{attn}}^{(l)})\|_{2}.$ (8)
For redundant tokens, such as those in homogeneous image regions, this error is negligible $(\epsilon^{(l)}\approx 0)$ due to minimal feature transformations by the FFN.
However, errors propagate through subsequent layers, amplified by the recursive nature of transformer architectures. Leveraging Lipschitz continuity assumptions for self-attention and FFN operations ($L^{(l+1)}_{attn}$ and $L^{(l+1)}_{FFN}$), the cumulative error at layer $l+1$ is bounded by:
$\epsilon^{(l+1)}\leq(L^{(l+1)}_{\text{attn}}+L^{(l+1)}_{\text{FFN}})\cdot\epsilon^{(l)}+\epsilon_{\text{skip}}^{(l+1)},$ (9)
where $(\epsilon_{\text{skip}}^{(l+1)}$ represents new errors from skipping deeper layers $(l+1)$.
Over $L$ layers, the total error telescopes to:
$\epsilon_{\text{total}}\leq\sum_{l=1}^{L}\epsilon_{\text{skip}}^{(l)}\cdot\prod_{i=1}^{L-l}(L^{(i+1)}_{\text{attn}}+L^{(i+1)}_{\text{FFN}}).$ (10)
Theorem 5.1
Bounded Lipschitz. If $W_{Q}$, $W_{K}$, $W_{V}$, $W_{1}$, $W_{2}$ are orthogonal matrices (spectral norm = 1), then:
  • $\text{L}(\text{attn})\leq 1/\sqrt{d_{k}}$
  • $\text{L}(\text{FFN})\leq 1$
Follow Theorem 5.1, we assume Lipschitz constants $L_{\text{attn}}+L_{\text{FFN}}\leq\gamma$ and skipping errors $\epsilon_{\text{skip}}^{(l)}\leq\epsilon$, the total error is scaled to:
$\epsilon_{\text{total}}\leq\epsilon\cdot\frac{\gamma^{L}-1}{\gamma-1},$ (11)
Equation 11 establishes that the skip error is bounded when $\gamma<1$, provided the model is trained with modern regularization techniques. This ensures that the MLLM remains less sensitive to the effects of skipping.
This error also impacts the KL divergence between the original and skipped outputs, bounded by:
$\mathcal{D}_{\text{KL}}(p_{\text{skip}}\parallel p_{\text{original}})\leq\frac{1}{2\sigma^{2}}\cdot\epsilon_{\text{total}}^{2},$ (12)
where $\sigma^{2}$ is the variance of the logits. Further integrating feature similarity errors $(\epsilon_{\text{sim}}=O(\sqrt{1-\theta}))$ from low-attention tokens, the final bound becomes:
$\mathcal{D}_{\text{KL}}\leq\frac{1}{2\sigma^{2}}\cdot\left(\epsilon_{\text{total}}+\epsilon_{\text{sim}}\right)^{2}.$ (13)
Practically, this analysis motivates a layer-wise skipping strategy, alongside token selection and merge based on feature similarity $(\theta)$. More analysis and the proof of Theorem 5.1 are provided in the Appendix.
6 Experiments
6.1 Benchmarks
To comprehensively evaluate our model, we conducted experiments on eight different benchmarks. These include MME [29], which assesses the perceptual and cognitive capabilities of multimodal language models; TextVQA [96], focused on fine-grained visual question answering; MMBench [72] and MMStar [20] for general diagnostic capabilities; MMMU [124], designed to test STEM-related reasoning; MathVista [76], which evaluates mathematical problem solving skills; OCRBench [70], for optical character recognition tasks; and MMVet [121], used for subjective assessment. Taken together, these benchmarks offer a comprehensive framework for assessing the efficacy of the model in various multimodal challenges.
6.2 Efficiency evaluation
We systematically evaluate both training time ($hours$), inference FLOP computation and inference latency ($ms$) on single A100 GPU. For training efficiency, we report actual GPU hours measured on consistent hardware configurations. For inference FLOPs, we adopt the methodology from FastV [21], specifically tracking the computational demands of processing visual tokens. In the LLama architecture, we calculate FLOPs for the causal attention and feed forward network modules as $4NC^{2}+2N^{2}C+3NCM$ where $N$ is the token count, $C$ denotes the hidden state dimension, and $M$ is the intermediate FFN dimension. Skip-Vision framework adjusts token counts dynamically between the attention and FFN modules, allowing us to refine FLOP calculations as $4N_{1}C^{2}+2N_{1}^{2}C+3N_{2}CM$, where $N_{1}$ and $N_{2}$ represent the visual token counts processed in the causal attention and FFN layers, respectively.
6.3 Data setup
During pretraining, we adopt the experimental settings of LLaVA and CoS, utilizing LLaVA-558K [69] and 65 million image-text pairs for respective comparisons. For the Supervised Fine-Tuning phase, we use LLaVA-665K [69] for all comparative experiments. By leveraging the computational efficiency of Skip-Vision framework, the resources saved can be redirected toward scaling the dataset to further enhance model performance. To this end, we extend our dataset to SV-1M, comprising 1 million samples, demonstrating the substantial performance improvements achievable through data expansion. To further assess our approach against state-of-the-art models of similar scale, we extend the dataset to SV-9M, underscoring the scalability and potential of our method. Detailed dataset statistics are provided in the Appendix.
6.4 Training setup
MME Textvqa MMB MMVet MMMU MathV OCRB MMStar Avg Hours FLOPs Latency
LLaVA setting:
LLaVA [69] 1535 58.6 72.2 32.8 39.6 21.3 32.7 36.5 42.0 10.2 100% 125
LLaVA-FastV [21] 1453 56.1 71 32.4 37.3 20.6 28.8 36.3 40.4 9.6 30% 103
MQT-LLaVA [39] 1487 53.6 69 28.2 37.8 16.8 28.1 36.5 38.6 7.5 44% 106
LLaVA-Tokenpacker [60] 1538 56.2 72.3 32.1 39.0 19.3 29.8 35.0 40.5 8.8 24% 97
SV-LLaVA 1519 56.5 72.5 32.4 38.8 21.2 30.7 36.5 41.3 7.9 25% 90
$(N_{r}=100,N_{s}=156)$ $-1\%$ $-3.6\%$ $+0.4\%$ $-1.2\%$ $-2\%$ $-0.5\%$ $-3.6\%$ $-6.1\%$ $-1.7\%$ $-22.5\%$ $-75\%$ $-28\%$
SV-LLaVA 1530 57.3 72.7 35.6 39.2 21.4 32.0 40.0 42.6 9.0 60% 103
$(N_{r}=256,N_{s}=316)$ $-0.3\%$ $-2.2\%$ $+0.7\%$ $+8.5\%$ $-1\%$ $+0.5\%$ $-2.1\%$ $+9.6\%$ $+1.4\%$ $-11.8\%$ $-40\%$ $-17.6\%$
LLaVA-HD setting:
LLaVA-HD [67] 1533 64 71.8 35.3 39.3 20 37.4 40.5 44.0 19 100% 450
SV-LLaVA-HD 1534 61.2 73.1 35.9 39.7 21.2 34.1 39.3 43.5 12.2 50% 250
$(N_{r}=576,N_{s}=576)$ $+0\%$ $-4.3\%$ $+1.8\%$ $+1.7\%$ $+1.5\%$ $-6\%$ $-8.8\%$ $-2.9\%$ $-1.1\%$ $-35.8\%$ $-50\%$ $-44.4\%$
CoS setting:
CoS [40] 1585 64.4 77.1 39.4 39.2 21.5 39.2 41.2 46.0 15.3 100% 160
SV-CoS 1563 63.7 76.5 41.7 40.3 21.2 37 41.9 46.0 9.8 25% 90
$(N_{r}=272,N_{s}=256)$ $-1.4\%$ $-1.1\%$ $-0.8\%$ $+5.8\%$ $+2.8\%$ $-1.4\%$ $-5.6\%$ $+1.7\%$ $-0\%$ $-40\%$ $-75\%$ $-43.8\%$
SV-CoS (SV-1M) 1569 67.1 77.1 43.3 40.7 22.0 50.7 44.4 49.3 16.3 25% 90
Table 2: Performance and efficiency evaluation. We evaluate Skip-Vision on LLaVA, LLaVA-HD, and CoS and compare it with state-of-the-art efficiency optimization models under the LLaVA setting. $N_{r}$ and $N_{s}$ denote the number of retained and skipped tokens, respectively.
LLM MME Textvqa MMB MMVet MMMU MathV OCRB MMStar
Open-source models in 8B tier:
mPLUG-Owl3 [118] Qwen2 - 69.0 77.6 40.1 - 65.0 - 40.1
Cambrian [101] LLaMA3 1547 71.7 75.9 - 42.7 49.0 62.4 -
LLaVA [69] LLaMA3 1535 58.6 72.2 32.8 39.6 21.3 32.7 36.5
Mini-Gemini-HD [61] LLaMA3 1606 70.2 72.7 - 37.3 37.0 47.7 -
LLaVA-Next [56] LLaMa3 1604 64.6 72.1 - 41.7 36.3 49.0 -
Ovis [77] LLaMA3 - - 77.4 - 44.7 40.8 - 49.5
MiniCPM-V2.5 [117] LLaMA3 - 76.6 77.6 - 45.8 54.3 72.5 -
SV-CoS (SV-9M) LLaMA3 1550 69.1 78.4 51.7 43 61.3 69.4 52.3
Table 3: MLLM evaluation. We present the performance of SV-CoS on SV-9M, comparing it against the current SOTA models of a similar scale. The highest scores are highlighted in bold, while the second-highest scores are indicated with underlines.
MME Textvqa MMB MMVet MMMU MathV OCRB MMStar Avg Latency
SV-LLaVA (w/o SK) 1519 56.7 72.5 31.4 38.8 21.2 30.9 36.5 41.1 103
SV-LLaVA 1519 56.5 72.5 32.4 38.8 21.2 30.7 36.5 41.3 90
$+0\%$ $-0.4\%$ $+0\%$ $+3.2\%$ $+0\%$ $+0\%$ $-0.6\%$ $+0\%$ $+0.5\%$ $-9.7\%$
SV-LLaVA-HD (w/o SK) 1533 61.6 73.0 35.8 39.9 21.2 34.5 39.7 43.7 310
SV-LLaVA-HD 1534 61.2 73.1 35.9 39.7 21.2 34.1 39.3 43.5 250
$+0\%$ $-0.7\%$ $+0.1\%$ $+0.3\%$ $-0.5\%$ $+0\%$ $-1.2\%$ $-1\%$ $-0.5\%$ $-19.4\%$
SV-LLaVA-CoS (w/o SK) 1562 63.9 76.5 40.2 40.2 21.2 37.2 41.9 45.9 126
SV-LLaVA-CoS 1563 63.7 76.5 41.7 40.3 21.2 37 41.9 46.0 90
$+0\%$ $-0.3\%$ $+0\%$ $+3.7\%$ $+0.2\%$ $+0\%$ $-0.5\%$ $+0\%$ $+0.2\%$ $-28.6\%$
Table 4: Ablation on skip KV-cache. We evaluate SK (using skip KV-cache during inference) of Skip-Vision on different settings.
Figure 6: Ablation study. We performed an ablation study on each component of the Skip-Vision. SF (skip FFN), TM (token merge).
Figure 7: Visualization of attention map in MMVet.
Figure 8: Visualization of attention map in TextVQA.
For LLaVA and LLaVA-HD setting, we use input resolutions of 336 and a dynamic resolution ranging from 336 to 672, respectively, with CLIP ViT-L/336px [87] as the vision encoder and LLaMA3 8B Instruct [28] as the backbone. For CoS setting, pretraining is conducted at 224 resolution, training only the multi-scale visual resampler with CLIP ViT-L/224px [87]. The model processes 80 visual tokens, including 16 global and 64 local tokens, and scales to 336 and 1296 tokens during SFT. In SFT, the resolution increases to 448, and all model parameters are unfrozen for full optimization.
In all experiments, we used a learning rate of $1e^{-3}$ for pre-training, which is reduced to $1e^{-5}$ during SFT. Training follows a cosine decay schedule after a $3\%$ warm-up phase. For skipped token selection, unless otherwise specified, the LLaVA setting includes 156 skipped tokens and 100 retained tokens, the LLaVA-HD setting has 576 skipped and 576 retained tokens, and the CoS setting consists of 256 skipped and 272 retained tokens. All skipped tokens undergo a token merging process to enhance efficiency.
6.5 Main experimental results
Performance and efficiency. We first evaluate the effectiveness of Skip-Vision across three different frameworks: CoS, LLaVA, and LLaVA-HD, demonstrating its ability to maximize efficiency while minimizing performance degradation. As shown in Figure 1 and Table 5, Skip-Vision reduces training time by $22.5\%$, inference FLOPs by $75\%$, and inference latency by $28\%$ for LLaVA. For LLaVA-HD, it achieves a $35.8\%$ reduction in training time, $50\%$ in inference FLOPs, and $45\%$ in inference latency. In CoS, it reduces training time by $35.8\%$, inference FLOPs by $75\%$, and inference latency by $44\%$.
Scaling. We evaluated the scalability of Skip-Vision, leveraging CoS’s visual scaling to increase visual tokens from 336 to 1296 while maintaining efficiency. Expanding the fine-tuning dataset from LLaVA-665k to SV-1M further improved performance by 8% over $CoS_{1296}$ with similar resources. To benchmark against state-of-the-art models, we scaled to SV-9M, with training completed in 72 hours on 16 NVIDIA A100 GPUs (Table 3).
6.6 Ablation study and analysis
To validate the effectiveness of each component within the Skip-Vision framework, we conducted ablation and comparative experiments on token merge, skip FFN, and skip KV-cache.
As shown in Figure 6, both the skip FFN strategy and token merging enhance computational efficiency across different model settings while maintaining performance. Notably, skip FFN performs better in the CoS setting than in LLaVA and LLaVA-HD. We attribute this to CoS allowing the vision backbone to be fine-tuned during the SFT stage, enabling key information to be more effectively captured by summary tokens, thereby reducing the errors introduced by skip FFN. As shown in Table 4, skip KV-cache slightly affects fine-grained tasks such as TextVQA and OCRBench but improves performance in long-form response tasks like MMVet.
To further analyze this, we visualized attention maps across layers for visual tokens in SV-CoS during the MMVet and TextVQA tasks. In MMVet (Figure 7), the attention of generated tokens primarily focuses on retained and summary tokens, with skip KV-cache filtering out unnecessary noise. In TextVQA (Figure 8), which requires fine-grained understanding, attention is spread across both retained and skipped tokens, leading to some performance loss with skip KV-cache. However, causal attention effectively aggregates information from skipped tokens into the final summary tokens, allowing us to retain only the summary tokens and retained visual tokens in the KV-cache.
7 Conclusion
We propose Skip-Vision, an efficient multimodal large language model architecture that boosts computational efficiency through a skip-FFN strategy, reducing redundant computations over visual tokens during training. During inference, Skip-Vision uses a skip KV-cache to accelerate processing by omitting non-essential visual tokens from the KV cache. Extensive experiments show our model outperforms current methods in efficiency and scales effectively with data, delivering competitive performance among leading models of similar scale.
Supplementary Material
8 Proof and detailed analysis of Skip-Vision
In this section, we provide a theoretical analysis to justify the rationale behind Skip-Vision and quantify the performance loss introduced by skipping FFN computations for redundant visual tokens.
8.1 Bounded error of Skip FFN
At the core of this analysis lies the layer error incurred when bypassing the FFN layer. For a given layer $l$, the original output $h_{\text{original}}^{(l)}$ is :
$h_{\text{original}}^{(l)}=h_{\text{attn}}^{(l)}+\text{FFN}^{(l)}(h_{\text{attn}}^{(l)}).$ (14)
When skipping the FFN, the output becomes:
$h_{\text{skip}}^{(l)}=h_{\text{attn}}^{(l)}.$ (15)
The per-layer skipping error is:
$\epsilon^{(l)}=\|h_{\text{original}}^{(l)}-h_{\text{skip}}^{(l)}\|_{2}=\|\text{FFN}^{(l)}(h_{\text{attn}}^{(l)})\|_{2}.$ (16)
For redundant tokens, such as those in homogeneous image regions, this error is negligible $(\epsilon^{(l)}\approx 0)$ due to minimal feature transformations by the FFN.
However, errors propagate through subsequent layers, amplified by the recursive nature of transformer architectures. Leveraging Lipschitz continuity assumptions for self-attention and FFN operations ($L^{(l+1)}_{attn}$ and $L^{(l+1)}_{FFN}$), the cumulative error at layer $l+1$ is bounded by:
$\epsilon^{(l+1)}\leq(L^{(l+1)}_{\text{attn}}+L^{(l+1)}_{\text{FFN}})\cdot\epsilon^{(l)}+\epsilon_{\text{skip}}^{(l+1)},$ (17)
where $(\epsilon_{\text{skip}}^{(l+1)}$ represents new errors from skipping deeper layers $(l+1)$.
Over $L$ layers, the total error telescopes to:
$\epsilon_{\text{total}}\leq\sum_{l=1}^{L}\epsilon_{\text{skip}}^{(l)}\cdot\prod_{i=1}^{L-l}(L^{(i+1)}_{\text{attn}}+L^{(i+1)}_{\text{FFN}}).$ (18)
Theorem 8.1
Lipschitz Constants for Causal Attention and FFN in Transformers Assume:
  1. 1.
    Inputs are normalized (e.g., via LayerNorm), bounding intermediate feature norms.
  2. 2.
    Weight matrices in attention ($W_{Q}$, $W_{K}$, $W_{V}$) and FFN ($W_{1}$, $W_{2}$) have bounded spectral norms (maximum singular values).
Then the Lipschitz constants of these components satisfy:
  1. 1.
    Causal Attention:
    $\text{L}(\text{Attn})\leq\frac{\|W_{Q}\|_{2}\|W_{K}\|_{2}\|W_{V}\|_{2}}{\sqrt{d_{k}}},$ (19)
    where $d_{k}$ is the key dimension. The causal mask further restricts attention dependencies, preserving this bound.
  2. 2.
    Feed - Forward Network (FFN):
    $\text{L}(\text{FFN})\leq\|W_{1}\|_{2}\|W_{2}\|_{2},$ (20)
    assuming the activation function (e.g., ReLU, GELU) is 1-Lipschitz.
Proof.
The Lipschitz constant of causal attention
1.Linear Transformation:
The input sequence $X\in\mathbb{R}^{n\times d}$ undergoes three linear transformations to obtain the query $Q=XW_{Q}$, the key $K=XW_{K}$, and the value $V=XW_{V}$. The Lipschitz constant of each linear transformation is the spectral norm (the largest singular value) of its weight matrix, denoted as $\sigma_{Q}=\|W_{Q}\|_{2}$, $\sigma_{K}=\|W_{K}\|_{2}$, and $\sigma_{V}=\|W_{V}\|_{2}$ respectively.
2.Attention Score Calculation:
The scaled dot - product $S=QK^{\top}/\sqrt{d_{k}}$. The Lipschitz constant of the bilinear mapping is related to $\sigma_{Q}$ and $\sigma_{K}$. If the norm of the input $X$ is bounded (for example, through LayerNorm), then:
$\text{Lip}(S)\leq\frac{\sigma_{Q}\sigma_{K}}{\sqrt{d_{k}}}.$ (21)
3. Softmax Activation:
After applying the causal mask, softmax is performed on each row. The Lipschitz constant of Softmax under the $\ell_{2}$ norm is less than 1, that is:
$\text{Lip}(\text{softmax})\leq 1.$ (22)
4. Weighted Sum of Values: The output $\text{Attn}(X)=AV$, where $A=\text{softmax}(S)$. The Lipschitz constant of this step is determined by the spectral norm $\sigma_{V}$ of the linear transformation of $V$.
Overall Lipschitz Constant
Combining the upper bounds of each step:
$\text{Lip}(\text{CausalAttention})\leq\frac{\sigma_{Q}\sigma_{K}\sigma_{V}}{\sqrt{d_{k}}}.$ (23)
Impact of Causal Mask: The mask restricts the attention range, which may reduce the sensitivity to the input. Therefore, the actual Lipschitz constant will not exceed the above-mentioned upper bound.
The Lipschitz Constant of FFN
The FFN is usually expressed as:
$\text{FFN}(x)=W_{2}\cdot\text{Activation}(W_{1}x+b_{1})+b_{2},$ (24)
where the Lipschitz constant of activation functions (such as ReLU, GELU) is 1.
Derivation of Lipschitz Constant
1. Linear Layer $W_{1}$: The spectral norm is $\sigma_{1}=\|W_{1}\|_{2}$.
2. Activation Function: $\text{Lip}(\text{Activation})=1$.
3. Linear Layer $W_{2}$: The spectral norm is $\sigma_{2}=\|W_{2}\|_{2}$.
Follow the equation 7 in [104], the overall Lipschitz constant is the product of the spectral norms of the two linear layers:
$\text{Lip}(\text{FFN})\leq\sigma_{1}\sigma_{2}.$ (25)
$\square$
Corollary 8.1
Bounded Lipschitz If $W_{Q}$, $W_{K}$, $W_{V}$, $W_{1}$, $W_{2}$ are orthogonal matrices (spectral norm = 1), then:
  • $\text{L}(\text{attn})\leq 1/\sqrt{d_{k}}$
  • $\text{L}(\text{FFN})\leq 1$
If we assume Lipschitz constants $L_{\text{attn}}+L_{\text{FFN}}\leq\gamma$ and skipping errors $\epsilon_{\text{skip}}^{(l)}\leq\epsilon$, the total error is scaled to:
$\epsilon_{\text{total}}\leq\epsilon\cdot\frac{\gamma^{L}-1}{\gamma-1},$ (26)
Theorem 5.1 establish that the skip error is bounded when $\gamma<1$, provided the model is trained with modern regularization techniques. This ensures that the Multimodal Large Language Model (MLLM) remains less sensitive to the effects of skipping.
This error also impacts the KL divergence between the original and skipped outputs, bounded by:
$\mathcal{D}_{\text{KL}}(p_{\text{skip}}\parallel p_{\text{original}})\leq\frac{1}{2\sigma^{2}}\cdot\epsilon_{\text{total}}^{2},$ (27)
where $\sigma^{2}$ is the variance of the logits.
Proof. For two Gaussian distributions $p=\mathcal{N}(\mu_{p},\Sigma_{p})$ and $q=\mathcal{N}(\mu_{q},\Sigma_{q})$, their KL divergence is:
$\mathcal{D}_{\text{KL}}(p\|q)$
$=\frac{1}{2}(\text{tr}(\Sigma_{q}^{-1}\Sigma_{p})+(\mu_{q}-\mu_{p})^{T}\Sigma_{q}^{-1}(\mu_{q}-\mu_{p})$
$-k+\ln\frac{|\Sigma_{q}|}{|\Sigma_{p}|})$ (28)
where $k$ is the dimension. If we assume that the covariances of the two distributions are the same, i.e., $\Sigma_{p}=\Sigma_{q}=\sigma^{2}I$, and the total difference in means is $\epsilon_{\text{total}}$, then:
$\mathcal{D}_{\text{KL}}(p\|q)=\frac{1}{2\sigma^{2}}\|\mu_{p}-\mu_{q}\|^{2}.$ (29)
Here, $\|\mu_{p}-\mu_{q}\|^{2}$ is $\epsilon_{\text{total}}^{2}$, so:
$\mathcal{D}_{\text{KL}}(p\|q)\leq\frac{1}{2\sigma^{2}}\epsilon_{\text{total}}^{2}.$ (30)
$\square$
Further integrating feature similarity errors $(\epsilon_{\text{sim}}=O(\sqrt{1-\theta}))$ from low-attention tokens, the final bound becomes:
$\mathcal{D}_{\text{KL}}\leq\frac{1}{2\sigma^{2}}\cdot\left(\epsilon_{\text{total}}+\epsilon_{\text{sim}}\right)^{2}.$ (31)
Practically, this analysis motivates a layer-wise skipping strategy, alongside token selection and token merge based on feature similarity $(\theta)$.
9 More experimental results
9.1 Efficiency.
Under the cos setting, we conduct more experiments to evaluate the training and inference efficiency. We compare with methods: FastV [21], Victor [109] and mean average pool, fine-tuning on LLaVA-665k using 8 NVIDIA A100 GPUs. As shown in Figure 11, our architecture outperforms in both metrics. Compared to the $CoS_{1296}$ baseline, it achieves comparable performance with $35\%$ less training time and $74\%$ reduced inference computation. FastV, unable to utilize flash attention, shows a significant disadvantage, even surpassing baseline training time.
Figure 9: Visualization of attention map in MME.
Figure 10: Visualization of attention map in TextVQA.
Figure 11: Performance vs. Training-Time and Inference FLOPs. Under the cos setting, we compare Skip-Vision with three MLLM acceleration methods, showing clear advantages in training speed and inference efficiency under equivalent computational constraints.
9.2 Ablation study
SF FS LS Merge LV SK MME Textvqa MMB MMVet MMMU MathV OCRB MMStar Overall
0 $CoS_{1296}$ 1585 64.4 77.1 39.4 39.2 21.5 39.2 41.2 46
1 $\checkmark$ 1548 64.5 75.3 39.7 39.1 21.8 37.7 40.9 45.6
2 $\checkmark$ 1589 63.6 74.4 36.9 38 19.8 365 39.8 44.2
3 $\checkmark$ $\checkmark$ 1591 63.7 74.9 39.4 38.7 20.7 36.3 39.9 44.8
4 $\checkmark$ $\checkmark$ 1560 63.8 75.5 39.0 39.2 21 37.1 40.0 45.0
5 $\checkmark$ $\checkmark$ $\checkmark$ 1580 63.5 74.2 40.6 39.2 21.6 37.2 40.9 45.3
6 $\checkmark$ $\checkmark$ $\checkmark$ 1570 63.5 74.1 40 39.8 20.1 39 41.4 45.4
7 $\checkmark$ $\checkmark$ $\checkmark$ 1593 62.6 74.7 40.3 40.2 21.6 36.6 41.1 45.3
8 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1571 63.6 75.9 40.3 40.3 21 36.1 41.5 45.6
9 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1547 64.0 75.9 38 41.2 21.9 36.9 40.6 45.5
10 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1562 63.9 76.5 40.2 40.2 21.2 37.2 41.9 45.9
11 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1563 63.7 76.5 41.7 40.3 21.2 37.0 41.9 46
Table 5: Ablation study. To establish a strong baseline, we performed an ablation study on each component of the Skip-Vision framework with the LLAVA 665k SFT dataset. SF (skip FFN), FS (former summary token), LS (latter summary token), Merge (reducing local visual tokens from 1024 to 256), LV (passing the last local visual token through the FFN), SK (using skip KV-cache during inference). This analysis highlights the distinct contributions of each element to efficiency and performance.
Inference method Skip window size MME Textvqa MMB MMVet MMMU MathV OCRB MMStar Overall
Without skip KV-cache - 1562 63.9 76.5 40.2 40.2 21.2 37.2 41.9 45.9
Skip KV-cache middle+small 1562 61.8 76.5 32.6 40.4 21.2 22.6 41.9 42.4
Skip KV-cache small 1563 63.7 76.5 41.7 40.3 21.2 37.0 41.9 46
Table 6: Ablation study of skip KV-cache.We report skip KV-cache performance across different visual token window sizes. Skip-Vision enables task-specific optimization by adjusting skip KV-cache levels for tailored acceleration.
To validate the effectiveness of each component within the Skip-Vision framework, under CoS setting, we conducted ablation and comparative experiments on the skip FFN, summary token, token merge, last visual token, and skip KV-cache. The detailed experimental results are presented in the table 5.
The last summary token or final visual token must pass through the FFN. As discussed in Section 4.2, the final visual token plays a crucial role in predicting the subsequent text token, thereby requiring access to the textual knowledge encoded within the FFN layers. This integration of information is essential. Compare (2, 3, 4, 5) in Table 5, employing a summary token as the final visual token has demonstrated enhanced effectiveness compared to a standard visual token. Furthermore, our experimental findings reveal that optimal performance is achieved when both the summary token and the last local visual token are processed through the FFN.
The former summary token enhances the model’s comprehension of overall visual information. Compare (4,6), (7,8), (9,10) in Table 5, the former summary token enhances the emphasis on crucial information by adaptively merging large-scale visual features. This approach addresses the challenge posed by overly lengthy sequences of visual tokens bypassing the FFN, which may result in the omission of critical large-scale visual context.
The local token merge strategy seamlessly aligns with the skip-vision framework. Compare (0,1), (4,7), (6,8) in Table 5, when loacl token merge is directly applied to the $CoS_{1296}$ baseline, the model’s overall performance declines, reflecting its dependency on redundant visual data when all tokens pass through the FFN. In contrast, within the skip-vision framework, merging local tokens results in improved performance, indicating that our architecture efficiently leverages visual information without requiring excessive redundancy.
The skip KV-cache mechanism enables adaptive selection of visual tokens to skip based on task-specific information requirements. As demonstrated in Table 6, for tasks such as MMB, MMU, and MMStar that do not necessitate fine-grained information, both middle and small window-size visual tokens can be skipped, with only the initial and final summary tokens retained. For detail-sensitive tasks like TextVQA and OCRBench, we skip only the small window-size tokens that bypass the FFN, thereby preserving critical fine-grained details. Applying skip KV-cache with small window-size tokens during inference improves performance, particularly in tasks requiring extended responses, such as MMVet.
9.3 More visualizations
Figure 12: Visualization of attention map in MMBench.
Figure 13: Visualization of attention map in MMVet.
Figure 14: Visualization of attention map in MMMU.
Figure 15: Visualization of attention map in MathV.
Figure 16: Visualization of attention map in OCRBench.
Figure 17: Visualization of attention map in MMStar.
9.4 Dataset
In Table 7 and Table 8, we introduce the two scaling datasets used by Skip-Vision during the SFT stage: SK-1M and SK-9M.
Task Dataset
Visual Instruction Tuning LLaVA-665k [69], SVIT [129]
VQA CREPE [122],Imagenet multi task [90], VQA-rad [52]
Visual Reasoning Wikitable [51], Super-CLEVR [62], VSR [65]
Knowledge ViQuAE [55],Kvqa [91], Websrc [22]
Chart / Diagram / graph ChartQA [79],Iconqa [74], Infographicvqa [82],
Document DeepForm [99], TAT-QA [132], Visualmrc [100], Docvqa [81], Sujet-Finance-QA-Vision-100k [98]
Math Mathverse [126]
OCR / Screen / Scene text TQA [50], HW-SQuAD [80], TextVQA [96], ST-VQA [8],TextOCR-GPT4V [13], OCRbench-kv [71], Uber-text [127]
Science AI2D [49]
Table 7: Datasets used by SK-1M at the SFT stage.
Task Dataset
Visual Instruction Tuning SVIT [129], ALLaVA [17], ShareGPT4V [19], cog-vlm-sft [106]
Caption TextCaps [95], ShareGPT-4o [107]
VQA CREPE [122],Imagenet multi task [90], VQA-rad [52], VQAv2 [4], Vizwiz [33]
Visual Reasoning Wikitable [51], Super-CLEVR [62], VSR [65], FigureQA [45], TallyQA [1], Visual cot [93], CLEVR [43], Raven [125]
Knowledge ViQuAE [55],Kvqa [91], Websrc [22], OK-VQA [78], Volcano [54], RLAIF-V [120]
Chart / Diagram / graph ChartQA [79],Iconqa [74], Infographicvqa [82], MapQA [16], TabFact [111], Chart2Text [46], DVQA [44], Chartbench [113], MMC [66]
Document DeepForm [99], TAT-QA [132], Visualmrc [100], Docvqa [81], Sujet-Finance-QA-Vision-100k [98], Docmatix [53], DocReason25K [37], DocStruct4M [38]
Math Mathverse [126], MathOCR [15], MathV360K [94], GeoGPT4V [11], Geo170K QA [30]
OCR / Screen / Scene text TQA [50], HW-SQuAD [80], TextVQA [96], ST-VQA [8],TextOCR-GPT4V [13], OCRbench-kv [71], Uber-text [127], OCR-VQA [86], ScreenQA [5], SynthText [32], ChromeWriting [110], K12 Printing [57], SQuAD [88], ICDAR19-LSVT [48], ICPR18-MTWI [35], ICDAR19-ArT [25], COCO-Text [103], Docscan [97], HierText [73]
Science AI2D [49], Plotqa [85], ArXivQA [59], ScienceQA [75]
Table 8: Datasets used by SK-9M at the SFT stage.